### v2

@@ Abstract
We consider Bayesian inference for image deblurring with total variation (TV) prior. Since the posterior is analytically intractable, we resort to Markov chain Monte Carlo (MCMC) methods. However, since most MCMC methods significantly deteriorate in high dimensions, they are not suitable to handle high resolution imaging problems. In this paper, we show how low-dimensional sampling can still be facilitated by exploiting the sparse conditional structure of the posterior. To this end, we make use of the local structures of the blurring operator and the TV prior by partitioning the image into rectangular blocks and employing a blocked Gibbs sampler with proposals stemming from the Metropolis-Hastings adjusted Langevin Algorithm (MALA). We prove that this MALA-within-Gibbs (MLwG) sampling algorithm has dimension-independent block acceptance rates and dimension-independent convergence rate. In order to apply the MALA proposals, we approximate the TV by a smoothed version, and show that the introduced approximation error is evenly distributed and dimension-independent. Since the posterior is a Gibbs density, we can use the Hammersley-Clifford Theorem to identify the posterior conditionals which only depend locally on the neighboring blocks. We outline computational strategies to evaluate the conditionals, which are the target densities in the Gibbs updates, locally and in parallel. In two numerical experiments, we validate the dimension-independent properties of the MLwG algorithm and demonstrate its superior performance over MALA.
@@ 1 Introduction
Image acquisition systems usually capture only a blurred version of the “true image”. In many applications, such as remote sensing, medical imaging, astronomy, and digital photography, the aim is to reconstruct the “true image” from the acquired image [3, 4]. Since the blurring can oftentimes be modeled by a linear convolution operator, image deblurring is a classical linear inverse problem. In this paper, we follow the typical assumptions that the blurred image is obtained by a linear operator, and that it is corrupted by additive Gaussian noise. Computing a solution is usually not straightforward due to the ill-posedness of the problem, which is due to the ill-conditioning of the forward operator and the noise. To resolve the ill-posedndess, we often require regularization, and a commonly used regularization technique in image reconstruction is the edge-preserving total variation regularization (TV) introduced in [30].
@@
(1) $\pi(\mathbf{x})\propto\exp(-l(\mathbf{x})-\varphi_{0}(\mathbf{x})),$
@@
In this paper, we show how low-dimensional sampling can nevertheless be facilitated by exploiting the sparse conditional structure of the posterior. The sparse conditional structure is due to the fact that the full conditional posterior of any image patch of arbitrary size only depends on the neighboring pixels within some radius. This can be intuitively understood by considering that the convolution as well as TV operate locally on the image. Mathematically, the independence relationships among image patches can be proven by the Hammersley-Clifford Theorem, because the considered posterior is a Gibbs density with a sparse neighborhood [18]. The Gibbs sampler is an attractive choice for posteriors with sparse conditional structure, because the reduced dependencies only require a local evaluation of the conditionals, which is usually cheaper to compute. Moreover, several updates may be performed in parallel. To make use of these advantages, we partition the image into square blocks of equal size and employ a blocked Gibbs sampler with local and parallel updates of the image blocks.
@@
(2) $\pi_{\varepsilon}(\mathbf{x})\propto\exp(-l(\mathbf{x})-\varphi_{\varepsilon}(\mathbf{x})).$
@@
It is shown in [32] that under the assumption of sparse conditional structure, MLwG can have dimension-independent block acceptance rate (at fixed step size) and dimension-independent convergence rate. In this paper, we present similar guarantees with specific conditions tailored to our target density and taking the approximation error into account. Finally, by making use of the sparse conditional structure of our posterior, which still holds under the smoothing, we show how an efficient and parallel MLwG algorithm can be implemented. We test the algorithm in two numerical experiments, where we illustrate the dimension-independent block acceptance rate and the dimension-independent convergence. Moreover, we compare the MLwG algorithm to MALA and show that MLwG clearly outperforms MALA in terms of sample quality and computational wall-clock time with increasing dimension.
@@
  • We develop a Stein’s method based approach to prove that the low dimensional marginals of the target distribution $\pi$ given in Eq. 1 and the smoothed distribution $\pi_{\varepsilon}$ in Eq. 2 have dimension-independent error in their marginals.
  • We show how to implement an efficient parallel MLwG sampling algorithm by providing the local target densities and their gradients for the block updates.
  • We illustrate the dimension-independent block acceptance rate and the dimension independent convergence in two numerical experiments, where we also show that MLwG clearly outperforms MALA.
@@
In this paper, we work with discrete and square images of $n\times n$ pixels, but the results can be extended to rectangular images. We employ a vector notation for the images by stacking them in the usual column-wise fashion. Furthermore, we partition the images into blocks of size $m\times m$, such that $n/m$ is an integer. We denote the number of pixels in one image by $d=n^{2}$, the number of pixels in one block by $q=m^{2}$, and the number of blocks by $b=(n/m)^{2}$. We denote pixels by lower case Greek letters, e.g., $\mathbf{x}_{\alpha}\in\mathbb{R}$, and blocks by lower case Roman letters, e.g., $\mathbf{x}_{i}\in\mathbb{R}^{q}$. We write $[b]\coloneqq\{1,2,\dots,d\}$ where $b$ is some positive integer, and $\setminus i\coloneqq[b]\setminus i$ for $i\in[b]$. We use upper case Greek letters to denote sets of pixel indices. For example, $\mathbf{x}_{\Theta}$ is the parameter block of pixels with indices in $\Theta\subseteq[d]$. We use the notation $\nabla_{i}:=\nabla_{\mathbf{x}_{i}}$ for the gradient operator with respect to the pixels of block $\mathbf{x}_{i}$.
@@
The remainder of this paper is organized as follows. In Section 2, we formulate the Bayesian deblurring problem and recall the blocked MLwG sampler. In Section 3, we propose the smoothing of the posterior and bound the error in the marginals in the Wasserstein-$1$ distance. In Section 4, we present the dimension-independent block acceptance rates and dimension-independent convergence rate of MLwG when applied to the smoothed posterior based on results from [32]. In Section 5, we present the local & parallel MLwG algorithm and show how the target densities, i.e., the conditionals, can be evaluated locally and in parallel. In Section 6, we validate the dimension-independent properties of the local & parallel MLwG algorithm in two numerical examples and perform a comparison to MALA. We conclude this paper with a short summary in Section 7.
@@ 2.1 Problem setting
We fist consider the classic image deconvolution problem with TV regularization, e.g., [24] and assume that a blurred and noisy image $\mathbf{y}\in\mathbb{R}^{d}$ is obtained by
(3) $\mathbf{y}=\mathbf{A}\mathbf{x}_{\mathrm{true}}+\mathbf{\epsilon}.$
Here, $\mathbf{x}_{\mathrm{true}}\in\mathbb{R}^{d}$ is the “true” image, $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\tfrac{1}{\lambda}\mathbf{I}_{d})$, and $\mathbf{A}\in\mathbb{R}^{d\times d}$ is the convolution operator. In particular, one can construct $\mathbf{A}$ via the discrete point spread function (PSF), i.e., the convolution kernel. We assume that the discrete PSF has radius $r>0$, such that $\mathbf{A}\mathbf{x}$ convolves each pixel with the surrounding $(2r+1)^{2}$ pixels.
Equation 3 constitutes an inverse problem where the goal is to recover a solution, which is “close to $\mathbf{x}_{\mathrm{true}}$” from the data $\mathbf{y}$. Computing a solution is typically not straightforward due to the ill-posedness of the problem. That is, $\mathbf{A}$ may be badly conditioned and thus highly sensitive to the noise. For this reason, we employ the edge-preserving TV regularization introduced in [30], which is a commonly used regularization technique in image reconstruction. For discretized images, it reads
(4) ${\rm TV}(\mathbf{x})=\sum_{\alpha=1}^{d}\sqrt{(\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}},$
where $\mathbf{D}^{(h)}\in\mathbb{R}^{d\times d}$ and $\mathbf{D}^{(h)}\in\mathbb{R}^{d\times d}$ are finite difference matrices for the computation of the horizontal and vertical differences between the pixels. The subscript $\alpha$ denotes the $\alpha$-th row of $\mathbf{D}^{(v)}$ and $\mathbf{D}^{(h)}$, such that $\mathbf{D}_{\alpha}^{(v)}\mathbf{x}$ and $\mathbf{D}_{\alpha}^{(h)}\mathbf{x}$ are the differences between pixel $\mathbf{x}_{\alpha}$ and its neighboring pixels in the vertical and horizontal directions, respectively. The finite difference matrices are defined as
@@
Here $\otimes$ denotes the Kronecker product. As for the convolution, we assume zero boundary conditions for the finite difference matrices. In practice, we compute the convolution and TV in matrix-free fashion. However, the matrix expressions are more convenient for the derivation of our theoretical results.
We now formulate the Bayesian inverse problem by defining the posterior probability density $\pi(\mathbf{x}|\mathbf{y})\propto\pi(\mathbf{y}|\mathbf{x})\pi_{\rm prior}(\mathbf{x})$, with the likelihood function $\mathbf{x}\mapsto\pi(\mathbf{y}|\mathbf{x})$ and the prior density $\pi_{\rm prior}(\mathbf{x})$. The likelihood function is determined by the data generating model Eq. 3, and reads
(6) $\pi(\mathbf{y}|\mathbf{x})\propto\exp\left(-\frac{\lambda}{2}\|\mathbf{y}-\mathbf{A}\mathbf{x}\|_{2}^{2}\right).$
We note that in this paper, the data $\mathbf{y}$ is fixed. The prior is constructed based on the discretized TV Eq. 4 and is of the Gibbs-type density
(7) $\pi_{\rm prior}(\mathbf{x})\propto\exp\left(-\delta{\rm TV}(\mathbf{x})\right),$
where $\delta>0$ is controlling the strength of the prior. In this paper, we consider $\delta$ to be given. The prior Eq. 7 is called the TV prior, and was introduced in [31, 13]. Consequently, the likelihood function Eq. 6 and the TV prior Eq. 7 give rise to a log-concave composite posterior density of the form
(8) $\begin{split}\pi(\mathbf{x}):=\pi(\mathbf{x}|\mathbf{y})&\propto\exp\left(-l(\mathbf{x})-\varphi_{0}(\mathbf{x})\right),\\ \text{with the potentials}\quad l(\mathbf{x})&:=\frac{\lambda}{2}\|\mathbf{y}-\mathbf{A}\mathbf{x}\|_{2}^{2}\quad\text{and}\quad\varphi_{0}(\mathbf{x}):=\delta{\rm TV}(\mathbf{x}).\end{split}$
@@ 2.2 MALA-within-Gibbs (MLwG)
A Gibbs sampler in its original form, also called component-wise or sequential Gibbs sampler, updates each component $\mathbf{x}_{\alpha}$ sequentially from the full conditional $\pi(\mathbf{x}_{\alpha}|\mathbf{x}_{\setminus\alpha})$, and a new sample $\mathbf{x}$ is obtained after all components are updated. It can be shown that a sample chain constructed by such an algorithm converges to the target density, see, e.g., [28]. Moreover, the convergence result holds also for block updates, i.e., blocked Gibbs sampling, which we employ in this paper.
To this end, we partition the image $\mathbf{x}$ into $b$ square block images $\mathbf{x}_{i}\in\mathbb{R}^{q}$, $i=1,\dots,b$. See Figure 1 for an example with a 4-by-4 block partition. The blocks $\mathbf{x}_{i}$ have equal side lengths $m\in\mathbb{N}$, such that $n/m$ is an integer, and they contain $q=m^{2}$ pixels. Moreover, we require $m>2r$, where $r$ is the radius of the discrete PSF in the convolution.
@@
We now establish some notation for the blocked Gibbs sampler. We call a complete iteration in which all blocks are updated a cycle, and denote by $\mathbf{x}^{k,i}\in\mathbb{R}^{d}$ the state during the $k$-th cycle after the update of the $i$-th block. To illustrate this notation, consider the following presentation of block updates:
(9) $\mathbf{x}^{0}\rightarrow\underbrace{\mathbf{x}^{1,1}\rightarrow\mathbf{x}^{1,2}\rightarrow\cdots\rightarrow\mathbf{x}^{1,b}}_{\text{$1$st cycle}}=:\mathbf{x}^{1}\rightarrow\underbrace{\mathbf{x}^{2,1}\rightarrow\mathbf{x}^{2,2}\rightarrow\cdots\rightarrow\mathbf{x}^{2,b}}_{\text{$2$nd cycle}}=:\mathbf{x}^{2}\rightarrow\cdots$
Notice also that we introduce $\mathbf{x}^{k}$ to denote the state after the $k$-th cycle. Other block updating rules like randomized sequences could also be employed, but are not discussed here.
Now consider the state $\mathbf{x}^{k,i-1}$ of the sample chain Eq. 9. To move to state $\mathbf{x}^{k,i}$, we fix $\mathbf{x}^{k,i}_{\setminus i}=\mathbf{x}^{k,i-1}_{\setminus i}$ and sample the remaining block $\mathbf{x}_{i}^{k,i}\sim\pi(\mathbf{x}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})$. If it is not possible to sample $\mathbf{x}_{i}^{k,i}\sim\pi(\mathbf{x}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})$ in closed form, one can use a single Metropolis-Hastings step as follows. A candidate $\mathbf{z}^{k,i}_{i}$ is simulated from a proposal density $q(\cdot|\mathbf{x}^{k,i-1}_{i})$ and accepted with probability
(10) $\alpha^{k,i}=1\wedge\frac{\pi(\mathbf{z}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})q(\mathbf{x}^{k,i-1}_{i}|\mathbf{z}^{k,i}_{i})}{\pi(\mathbf{x}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})q(\mathbf{z}^{k,i}_{i}|\mathbf{x}^{k,i-1}_{i})}.$
In MLwG, the MALA proposal density is used. For the target density $\pi(\mathbf{x}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})$, it reads
(11) $q(\mathbf{z}_{i}^{k,i}|\mathbf{x}_{i}^{k,i})=\exp\left(-\frac{1}{4\tau}\|\mathbf{z}_{i}^{k,i}-\mathbf{x}_{i}^{k,i}-\tau\nabla_{i}\log\pi(\mathbf{x}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})\|_{2}^{2}\right),$
@@
(12) $\mathbf{z}_{i}^{k,i}=\mathbf{x}_{i}^{k,i}+\tau\nabla_{i}\log\pi(\mathbf{x}_{i}^{k,i}|\mathbf{x}^{k,i-1}_{\setminus i})+\sqrt{2\tau}\mathbf{\xi}_{i}^{k,i},\quad\mathbf{\xi}_{i}^{k,i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{q}).$
@@
0:  Number of samples $N$, step size $\tau$, initial state $\mathbf{x}^{0}$.
@@
4:        Set $\mathbf{x}^{k,i}\leftarrow\mathbf{x}^{k,i-1}$.
5:        Propose block candidate $\mathbf{z}^{k,i}_{i}$ according to Eq. 12.

### v3

@@ Abstract
We consider Bayesian inference for image deblurring with total variation (TV) prior. Since the posterior is analytically intractable, we resort to Markov chain Monte Carlo (MCMC) methods. However, since most MCMC methods significantly deteriorate in high dimensions, they are not suitable to handle high resolution imaging problems. In this paper, we show how low-dimensional sampling can still be facilitated by exploiting the sparse conditional structure of the posterior. To this end, we make use of the local structures of the blurring operator and the TV prior by partitioning the image into rectangular blocks and employing a blocked Gibbs sampler with proposals stemming from the Metropolis-adjusted Langevin Algorithm (MALA). We prove that this MALA-within-Gibbs (MLwG) sampling algorithm has dimension-independent block acceptance rates and dimension-independent convergence rate. In order to apply the MALA proposals, we approximate the TV by a smoothed version and show that the introduced approximation error is evenly distributed and dimension-independent. Since the posterior is a Gibbs density, we can use the Hammersley-Clifford Theorem to identify the posterior conditionals which only depend locally on the neighboring blocks. We outline computational strategies to evaluate the conditionals, which are the target densities in the Gibbs updates, locally and in parallel. In two numerical experiments, we validate the dimension-independent properties of the MLwG algorithm and demonstrate its superior performance over MALA.
@@ 1 Introduction
Image acquisition systems usually capture only a blurred version of the “true image”. In many applications, such as remote sensing, medical imaging, astronomy, and digital photography, the aim is to reconstruct the “true image” from the acquired image [3, 4]. Since the blurring can oftentimes be modeled by a linear convolution operator, image deblurring is a classical linear inverse problem. In this paper, we follow the typical assumptions that the blurred image is obtained by a linear operator and that it is corrupted by additive Gaussian noise. Computing a solution is usually not straightforward due to the ill-posedness of the problem, which is due to the ill-conditioning of the forward operator and the noise. To resolve the ill-posedndess, we often require regularization, and a commonly used regularization technique in image reconstruction is the edge-preserving total variation regularization (TV) introduced in [32].
@@
(1) $\pi({\bm{x}})\propto\exp(-l({\bm{x}})-\varphi_{0}({\bm{x}})),$
@@
In this paper, we show how low-dimensional sampling can nevertheless be facilitated by exploiting the sparse conditional structure of the posterior. The sparse conditional structure is due to the fact that the full conditional posterior of any image patch of arbitrary size only depends on the neighboring pixels within some radius. This can be intuitively understood by considering that the convolution as well as TV operate locally on the image. Mathematically, the independence relationships among image patches can be proven by the Hammersley-Clifford Theorem because the considered posterior is a Gibbs density with a sparse neighborhood [19]. The Gibbs sampler is an attractive choice for posteriors with sparse conditional structure because the reduced dependencies only require a local evaluation of the conditionals, which is usually cheaper to compute. Moreover, several updates may be performed in parallel. To make use of these advantages, we partition the image into square blocks of equal size and employ a blocked Gibbs sampler with local and parallel updates of the image blocks.
@@
(2) $\pi_{\varepsilon}({\bm{x}})\propto\exp(-l({\bm{x}})-\varphi_{\varepsilon}({\bm{x}})).$
@@
It is shown in [34] that under the assumption of sparse conditional structure, MLwG can have dimension-independent block acceptance rate (at fixed step size) and dimension-independent convergence rate. In this paper, we present similar guarantees with specific conditions tailored to our target density and taking the approximation error into account. Finally, by making use of the sparse conditional structure of our posterior, which still holds under the smoothing, we show how an efficient local & parallel MLwG algorithm can be implemented. We test the algorithm in two numerical experiments, where we illustrate the dimension-independent block acceptance rate and the dimension-independent convergence. Moreover, we compare the MLwG algorithm to MALA and show that MLwG clearly outperforms MALA in terms of sample quality and computational wall-clock time with increasing dimension.
@@
  • We develop a Stein’s method based approach to prove that the error between the low dimensional marginals of the target distribution $\pi$ in Eq. 1 and the smoothed distribution $\pi_{\varepsilon}$ in Eq. 2 is dimension-independent.
  • We show how to implement an efficient local & parallel MLwG sampling algorithm by providing the local target densities and their gradients for the block updates.
  • We illustrate the dimension-independent block acceptance and convergence rates in two numerical experiments, where we also show that MLwG clearly outperforms MALA.
@@
In this paper, we work with discrete and square images of $n\times n$ pixels, but the results can be extended to rectangular images. We employ a vector notation for the images by stacking them in the usual column-wise fashion. Furthermore, we partition the images into blocks of size $m\times m$, such that $n/m$ is an integer. We denote the number of pixels in one image by $d=n^{2}$, the number of pixels in one block by $q=m^{2}$, and the number of blocks by $b=(n/m)^{2}$. We denote pixels by lower case Greek letters and blocks by lower case Roman letters. E.g., for an image ${\bm{x}}\in\mathbb{R}^{d}$, we write ${\bm{x}}_{\alpha}\in\mathbb{R}$, and ${\bm{x}}_{i}\in\mathbb{R}^{q}$. We write $[b]\coloneqq\{1,2,\dots,b\}$ where $b$ is some positive integer, and $\setminus i\coloneqq[b]\setminus i$ for $i\in[b]$. We use upper case Greek letters to denote sets of pixel indices. For example, ${\bm{x}}_{\Theta}$ is the parameter block of pixels with indices in $\Theta\subseteq[d]$. We use the notation $\nabla_{i}:=\nabla_{{\bm{x}}_{i}}$ for the gradient operator with respect to the pixels of block ${\bm{x}}_{i}$.
@@
The remainder of this paper is organized as follows. In Section 2, we formulate the Bayesian deblurring problem and recall the blocked MLwG sampler. In Section 3, we propose the smoothing of the posterior and bound the error in the marginals in the Wasserstein-$1$ distance. In Section 4, we present the dimension-independent block acceptance rates and dimension-independent convergence rate of MLwG when applied to the smoothed posterior based on results from [34]. In Section 5, we present the local & parallel MLwG algorithm and show how the target densities, i.e., the conditionals, can be evaluated locally and in parallel. In Section 6, we validate the dimension-independent properties of the local & parallel MLwG algorithm in two numerical examples and perform a comparison to MALA. We conclude this paper in Section 7.
@@ 2.1 Problem setting
We fist consider the classic image deconvolution problem with TV regularization, e.g., [25] and assume that a blurred and noisy image ${\bm{y}}\in\mathbb{R}^{d}$ is obtained by
(3) ${\bm{y}}={\bm{A}}{\bm{x}}_{\mathrm{true}}+\bm{\epsilon}.$
Here, ${\bm{x}}_{\mathrm{true}}\in\mathbb{R}^{d}$ is the “true” image, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\tfrac{1}{\lambda}\bm{I}_{d})$, and ${\bm{A}}\in\mathbb{R}^{d\times d}$ is the convolution operator. In particular, one can construct ${\bm{A}}$ via the discrete point spread function (PSF), i.e., the convolution kernel. We assume that the discrete PSF has radius $r>0$, such that ${\bm{A}}{\bm{x}}$ convolves each pixel with the surrounding $(2r+1)^{2}$ pixels.
Equation 3 constitutes an inverse problem where the goal is to recover a solution that is “close to ${\bm{x}}_{\mathrm{true}}$” from the data ${\bm{y}}$. Computing a solution is typically not straightforward due to the ill-posedness of the problem. That is, ${\bm{A}}$ may be badly conditioned and thus highly sensitive to the noise. For this reason, we employ the edge-preserving TV regularization introduced in [32], which is a commonly used regularization technique in image reconstruction. For discretized images, it reads
(4) ${\rm TV}({\bm{x}})=\sum_{\alpha=1}^{d}\sqrt{({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}},$
where ${\bm{D}}^{(v)}\in\mathbb{R}^{d\times d}$ and ${\bm{D}}^{(h)}\in\mathbb{R}^{d\times d}$ are finite difference matrices for the computation of the horizontal and vertical differences between the pixels. The subscript $\alpha$ denotes the $\alpha$-th row of ${\bm{D}}^{(v)}$ and ${\bm{D}}^{(h)}$, such that ${\bm{D}}_{\alpha}^{(v)}{\bm{x}}$ and ${\bm{D}}_{\alpha}^{(h)}{\bm{x}}$ are the differences between pixel ${\bm{x}}_{\alpha}$ and its neighboring pixels in the vertical and horizontal directions, respectively. The finite difference matrices are defined as
@@
Here, $\otimes$ denotes the Kronecker product. In this paper, without loss of generality, we consider zero boundary conditions for the finite difference matrices. Other boundary conditions can be used by slight modification of the algorithm and the analysis. In practice, we compute the convolution and TV in a matrix-free way due to the high dimensionality of the problems. However, the matrix expressions are more convenient for the derivation of our theoretical results.
We now formulate the Bayesian inverse problem by defining the posterior probability density $\pi({\bm{x}}|{\bm{y}})\propto\pi({\bm{y}}|{\bm{x}})\pi_{\rm prior}({\bm{x}})$ with the likelihood function ${\bm{x}}\mapsto\pi({\bm{y}}|{\bm{x}})$ and the prior density $\pi_{\rm prior}({\bm{x}})$. The likelihood function is determined by the data generating model Eq. 3 and reads
(6) $\pi({\bm{y}}|{\bm{x}})\propto\exp\left(-\frac{\lambda}{2}\|{\bm{y}}-{\bm{A}}{\bm{x}}\|_{2}^{2}\right).$
We note that the data ${\bm{y}}$ is fixed in this paper. The prior is constructed based on the discretized TV Eq. 4 and is of the Gibbs-type density
(7) $\pi_{\rm prior}({\bm{x}})\propto\exp\left(-\delta{\rm TV}({\bm{x}})\right),$
where $\delta>0$ is controlling the strength of the prior. In this paper, we consider $\delta$ to be given. The prior Eq. 7 is called the TV prior and was introduced in [33, 14]. Consequently, the likelihood function Eq. 6 and the TV prior Eq. 7 give rise to a log-concave composite posterior density of the form
(8) $\begin{split}\pi({\bm{x}}):=\pi({\bm{x}}|{\bm{y}})&\propto\exp\left(-l({\bm{x}})-\varphi_{0}({\bm{x}})\right),\\ \text{with the potentials}\quad l({\bm{x}})&:=\frac{\lambda}{2}\|{\bm{y}}-{\bm{A}}{\bm{x}}\|_{2}^{2}\quad\text{and}\quad\varphi_{0}({\bm{x}}):=\delta{\rm TV}({\bm{x}}).\end{split}$
@@ 2.2 MALA-within-Gibbs (MLwG)
A Gibbs sampler in its original form, also called component-wise or sequential Gibbs sampler, updates each component ${\bm{x}}_{\alpha}$ sequentially from the full conditional $\pi({\bm{x}}_{\alpha}|{\bm{x}}_{\setminus\alpha})$, and a new sample ${\bm{x}}$ is obtained after all components are updated. It can be shown that a sample chain constructed by such an algorithm converges to the target density, see, e.g., [29]. Moreover, the convergence result holds also for block updates, i.e., blocked Gibbs sampling, which we employ in this paper.
To this end, we partition the image ${\bm{x}}$ into $b$ square block images ${\bm{x}}_{i}\in\mathbb{R}^{q}$, $i=1,\dots,b$. See Figure 1 for an example with a 4-by-4 block partition. The blocks ${\bm{x}}_{i}$ have equal side lengths $m\in\mathbb{N}$, such that $n/m$ is an integer, and they contain $q=m^{2}$ pixels. Moreover, we require $m>2r$, where $r$ is the radius of the discrete PSF in the convolution.
@@
We now establish some notation for the blocked Gibbs sampler. We call a complete iteration in which all blocks are updated a cycle and denote by ${\bm{x}}^{k,i}\in\mathbb{R}^{d}$ the state during the $k$-th cycle after the update of the $i$-th block. To illustrate this notation, consider the following presentation of block updates:
(9) ${\bm{x}}^{0}\rightarrow\underbrace{{\bm{x}}^{1,1}\rightarrow{\bm{x}}^{1,2}\rightarrow\cdots\rightarrow{\bm{x}}^{1,b}}_{\text{$1$st cycle}}=:{\bm{x}}^{1}\rightarrow\underbrace{{\bm{x}}^{2,1}\rightarrow{\bm{x}}^{2,2}\rightarrow\cdots\rightarrow{\bm{x}}^{2,b}}_{\text{$2$nd cycle}}=:{\bm{x}}^{2}\rightarrow\cdots$
Notice also that we introduce ${\bm{x}}^{k}$ to denote the state after the $k$-th cycle. Other block updating rules like randomized sequences could also be employed, but are not discussed here.
Now consider the state ${\bm{x}}^{k,i-1}$ of the sample chain Eq. 9. To move to state ${\bm{x}}^{k,i}$, we fix ${\bm{x}}^{k,i}_{\setminus i}={\bm{x}}^{k,i-1}_{\setminus i}$ and sample the remaining block ${\bm{x}}_{i}^{k,i}\sim\pi({\bm{x}}_{i}^{k,i}|{\bm{x}}^{k,i-1}_{\setminus i})$. If it is not possible to sample ${\bm{x}}_{i}^{k,i}\sim\pi({\bm{x}}_{i}^{k,i}|{\bm{x}}^{k,i-1}_{\setminus i})$ in closed form, one can use a single Metropolis-Hastings step as follows. A candidate ${\bm{z}}^{k,i}_{i}$ is simulated from a proposal density $q(\cdot|{\bm{x}}^{k,i-1}_{i})$ and accepted with probability
(10) $\alpha^{k,i}=1\wedge\frac{\pi({\bm{z}}_{i}^{k,i}|{\bm{x}}^{k,i-1}_{\setminus i})q({\bm{x}}^{k,i-1}_{i}|{\bm{z}}^{k,i}_{i})}{\pi({\bm{x}}_{i}^{k,i-1}|{\bm{x}}^{k,i-1}_{\setminus i})q({\bm{z}}^{k,i}_{i}|{\bm{x}}^{k,i-1}_{i})}.$
Note that the proposal density $q(\cdot|{\bm{x}}^{k,i-1}_{i})$ depends on ${\bm{x}}^{k,i-1}_{\setminus i}$, but here we omit the dependency for clearness. In MLwG, the MALA proposal density is used. For the target density $\pi({\bm{x}}_{i}^{k,i}|{\bm{x}}^{k,i-1}_{\setminus i})$, it reads
(11) $q({\bm{z}}_{i}^{k,i}|{\bm{x}}_{i}^{k,i})=\exp\left(-\frac{1}{4\tau}\|{\bm{z}}_{i}^{k,i}-{\bm{x}}_{i}^{k,i}-\tau\nabla_{i}\log\pi({\bm{x}}_{i}^{k,i}|{\bm{x}}^{k,i-1}_{\setminus i})\|_{2}^{2}\right),$
@@
(12) ${\bm{z}}_{i}^{k,i}={\bm{x}}_{i}^{k,i}+\tau\nabla_{i}\log\pi({\bm{x}}_{i}^{k,i}|{\bm{x}}^{k,i-1}_{\setminus i})+\sqrt{2\tau}\bm{\xi}_{i}^{k,i},\quad\bm{\xi}_{i}^{k,i}\sim\mathcal{N}(\bm{0},\bm{I}_{q}).$
Here, $\tau>0$ is the step size or scale parameter. The choice of the step size is crucial for the performance of MALA, as too large step sizes lead to many rejected samples, and too small step sizes lead to samples which are very close to each other. In both cases, the correlation between the samples becomes large which is not preferable. Several adaptive strategies for the tuning of the step size exist, and in our numerical experiments in Section 6, we use the method from [20] which adapts the step size during burn-in. In the MLwG sampler, the adaptive strategies can be used to tune the step size for each block individually and automatically such that the algorithm requires little parameter tuning in general.
@@
0:  Number of samples $N$, step size $\tau$, initial state ${\bm{x}}^{0}$.
@@
4:        Set ${\bm{x}}^{k,i}\leftarrow{\bm{x}}^{k,i-1}$.
5:        Propose block candidate ${\bm{z}}^{k,i}_{i}$ according to Eq. 12.